### v1

Abstract
With the release of ChatGPT, OpenAI has reinforced its leadership in large language models. Simultaneously, OpenAI offers developers advanced technologies and resources (such as API services), driving innovation and development in various applications. This initiative promotes technological advancement and inspires developers to further explore and address the challenges of applying these technologies.
In contrast to traditional software development practices, OpenAI’s development introduces new challenges for AI developers in design, implementation, and deployment. These challenges span different areas (such as prompts, plugins, and APIs), requiring developers to navigate unique methodologies and considerations specific to large language model development.
However, to the best of our knowledge, these challenges have not been thoroughly investigated in previous empirical studies. To fill this gap, we present the first comprehensive study on understanding the challenges faced by OpenAI developers. Specifically, we crawl and analyze 29,057 relevant questions from a popular OpenAI developer forum. We first examine their popularity and difficulty.
After manually analyzing 2,364 sampled questions, we construct a taxonomy of challenges faced by OpenAI developers. Based on this taxonomy, we summarize a set of findings and actionable implications for OpenAI-related stakeholders, including developers and the OpenAI organization.
Keywords:
Mining Software Repository, Empirical Study, OpenAI Developer, Development Challenges, Prompt Engineering
journal: Empirical Software Engineering
1 Introduction
The emerging large language models (LLMs) are increasingly attracting attention and becoming one of the hot research topics in the field of computer science. Until now, LLMs have demonstrated promising performance in different fields, such as natural language processing and software engineering (Zhao et al, 2023b; Wang et al, 2024; Fan et al, 2023).
OpenAI, an artificial intelligence research organization, has been at the forefront of developing large language models (LLMs), which are AI systems designed to understand and generate human-like text. These models, such as GPT-3 and its successor GPT-4, have garnered significant attention for their ability to perform a wide range of language-based tasks with impressive performance.
With the rise in popularity of ChatGPT and the rapid development of related OpenAI technologies, an increasing number of developers are utilizing these tools to assist in their development processes.
However, during OpenAI development, developers often encounter various challenges. For instance, correctly configuring and invoking OpenAI’s API can be difficult, including setting parameters, managing rate limits, and handling errors. For those unfamiliar with AI and LLMs, developing plugins and applications based on OpenAI’s technology can be daunting, involving integration, performance optimization, and ensuring security. Ensuring data privacy and security while handling user data is crucial. Developers must comply with relevant regulations and implement necessary security measures.
Compared to traditional software development, OpenAI developers face the following unique challenges: (1) Automating task processing: OpenAI’s large language models can automatically handle many tasks, such as text generation, image recognition, and speech recognition. In contrast, traditional software engineering often requires manual coding of various algorithms and logic to accomplish these tasks. (2) Dealing with uncertainty: LLMs can produce variable and sometimes unpredictable outputs, unlike traditional software which typically has deterministic outcomes. Developers must account for this uncertainty and implement mechanisms to manage it. (3) Handling large-scale datasets: OpenAI development often involves working with large-scale datasets to train or fine-tune models. This requires specialized knowledge in data preprocessing, management, and the efficient use of computational resources. (4) Data privacy and security: Using LLMs requires a significant amount of data for training and fine-tuning, raising concerns about the privacy and security of user data. Developers need to design effective methods to ensure data security and privacy. (5) Performance optimization: Optimizing the performance of large language models (LLMs), such as output accuracy, differs from optimizing traditional software. (6) Interpreting model outputs: Understanding and interpreting the outputs of LLMs can be complex. Developers need to ensure that the model’s outputs are reliable and contextually appropriate. These challenges require a different set of skills and considerations, highlighting the distinct nature of OpenAI development compared to traditional software development. However, to the best of our knowledge, these challenges faced by OpenAI developers have not been thoroughly investigated in previous empirical studies.
To fill this gap, we present the first comprehensive study on understanding the challenges for OpenAI developers. Our empirical study aims to help developers to avoid common pitfalls and improve development efficiency. To this end, we crawl and analyze 29,057 questions from the OpenAI developer forum [^1] https://community.openai.com.
The OpenAI Developer Forum is a collaborative space for developers to seek assistance and share insights on using OpenAI technologies. It supports developers of all skill levels, offering discussions on API integration, plugin development, and best practices. The forum also provides updates on the latest advancements and tools from OpenAI.
In our empirical study, we want to answer the following three research questions.
RQ1: What is the popularity trend of OpenAI development among developers?
Result. After analyzing posts related to OpenAI development, we examine the number of new posts and new users added over each period. Our findings indicate that OpenAI development is attracting increasing attention from developers, particularly with the introduction of pivotal products like ChatGPT. This growing trend underscores the rising popularity of OpenAI technologies, highlighting the timeliness and importance of our empirical study.
RQ2: How difficult is OpenAI development for developers?
Result. To illustrate the difficulty level of the challenges faced by OpenAI developers, we analyze the number of replies to OpenAI development-related questions. The results show that 53% of these questions have fewer than three replies, suggesting that they are typically challenging to address. This finding prompts us to further investigate the underlying challenges of these development questions.
RQ3: What specific challenges for OpenAI developers?
Result. We perform a manual analysis on 2,364 sampled questions and construct a taxonomy of challenges consisting of 27 categories. For example, Prompt Design, Integration with Custom Applications, and Token Limitation. In addition, based on this taxonomy, we summarize findings and actionable implications for stakeholders (such as developers and the OpenAI organization).
The main contributions of our empirical study can be summarized as follows:
  • We conduct the first empirical study to investigate the challenges faced by OpenAI developers by mining the OpenAI developer forum.
  • Our study constructs a taxonomy with detailed findings on these challenges.
  • We share our collected posts, analysis scripts, and taxonomy of challenges in a GitHub repository for open science [^2] https://github.com/judeomg/OADF.
Paper Organization. The rest of this paper is organized as follows. Section 2 introduces the background of OpenAI’s development. Section 3 describes our methods for collecting and analyzing forum posts. Section 4 analyzes the popularity of the forum. Section 5 analyzes the difficulty levels of the posts in the forum. Section 6 details our constructed taxonomy and discusses related insights. Section 7 describes validity threats. Section 8 discusses work related to development challenges. Section 9 summarizes this research.
2 Background
OpenAI has emerged as a leading organization in artificial intelligence research, offering a range of APIs that enable developers to integrate advanced AI capabilities into their applications. The OpenAI API provides access to various models that can perform a wide array of tasks, from natural language understanding and generation to more complex functions like code completion and image recognition. This section provides an overview of the API usage and the development of plugins within the OpenAI ecosystem.
OpenAI Usage. The OpenAI API allows developers to leverage the power of state-of-the-art AI models through a simple interface. Key features of the API include: (1) Ease of Integration: Developers can integrate the API into their applications with minimal setup, using familiar HTTP requests. (2) Versatile Capabilities: The API supports various tasks including text generation, translation, summarization, and more. (3) Scalability: The API is designed to handle a wide range of requests, from small-scale personal projects to large-scale enterprise applications. The API offers several endpoints corresponding to different models and capabilities. For instance, the davinci model is known for its high performance in generating coherent and contextually relevant text, making it suitable for applications requiring sophisticated language understanding.
Plugin Development. Plugins are extensions that enhance the functionality of software applications by embedding AI capabilities directly within them. OpenAI provides robust support for plugin development, allowing developers to create custom plugins that interact with OpenAI models. Key aspects of plugin development include: (1) API Integration: Plugins typically make use of the OpenAI API to fetch responses from AI models based on user input or other triggers. (2) Custom Functionality: Developers can tailor the behavior of plugins to meet specific needs, such as automating customer support responses, generating content, or providing real-time data analysis. (3) Interoperability: OpenAI plugins can be integrated with various platforms and services, enhancing the AI capabilities of a wide range of applications from web and mobile apps to enterprise systems.
OpenAI has introduced a customized version of ChatGPT called GPTs [^3] https://openai.com/index/introducing-gpts. Similar to plugins, GPTs serve personalized user needs. Developers can create different GPTs for specific purposes, which can be shared with others. GPTs offer a new way to create a tailored version of ChatGPT, assisting in daily life and specific tasks. For instance, GPTs can assist in analyzing math problems, learning soccer rules, or even helping in designing icons. Moreover, creating GPTs is a simple process that does not require coding, allowing anyone to easily build their GPTs by engaging in conversation with ChatGPT, providing instructions and additional knowledge, and selecting actions like web searching, data analysis, or image creation.
While the integration of OpenAI’s APIs and the development of plugins and GPTs offer significant advantages, they also present several challenges, such as: (1) Cost Management: Managing the computational and financial costs associated with training, fine-tuning, and deploying large AI models, which can be significantly higher than those for traditional software systems. (2) Ethical Considerations: Ensuring that AI usage adheres to ethical guidelines, particularly in areas like content generation and decision-making. (3) Output authenticity: Verifying the accuracy and reliability of the output generated by artificial intelligence models to ensure that plugins and GPTs meet development goals and user expectations.
3 Methodology
To understand the challenges faced by OpenAI developers, we collect and analyze relevant question posts from the OpenAI developer forum, following previous studies (Ahmed and Bagherzadeh, 2018; Alshangiti et al, 2019; Bagherzadeh and Khatchadourian, 2019; Lou et al, 2020; Rosen and Shihab, 2016; Yang et al, 2016; Zhang et al, 2019). Fig. 1 provides a methodology overview of our empirical study.
Figure 1: Methodology overview of our empirical study.
Step 1: Crawling question posts. To analyze the popularity trends and difficulty levels of the questions, we crawl posts from the OpenAI developer forum. To analyze challenges faced by OpenAI developers, we resort to categories, which make it easier to find relevant posts and ensure that questions reach the appropriate experts. Our study focuses on four specific categories (i.e., API, Prompting, GPT builders, and ChatGPT). Specifically, posts in the “API” category concern questions, feedback, and best practices around building with OpenAI’s API. posts in the “Prompting” category learn more about prompting by sharing best practices, and favorite prompts. GPTs are a new way for anyone to create a tailored version of ChatGPT to be more helpful in their daily life, therefore, the “GPT builders” category is designed for people who are building GPTs. Finally, the “ChatGPT” concerns questions or discussions about ChatGPT. since posts in the remaining categories (i.e., Community, Announcements, Documentation, and Forum feedback) rarely discuss the challenges faced by OpenAI developers, we ignore these posts in our empirical study. Specifically, posts in the “Announcements” category typically contain official product updates, those in the “Documentation” category offer tutorials and instructional content, posts in the “Community” category involve project and technology sharing, and posts in the “Forum feedback” category focus on improving the developer forum.
Fig. 2 shows a screenshot of a post [^4] https://community.openai.com/t/547685 from the OpenAI developer forum. Typically, a post consists of several key elements: title, category, tags, question description, and code snippet, with a few posts having accepted answers. These posts also include additional relevant information, such as creation time, last reply time, reply count, view count, and the number of participating users. In this step, we gather the following metadata: the post’s title, creation date, and the number of replies. We also collect user metadata including usernames and registration times. In summary, by June 2024, we had crawled data from 29,057 posts and 722,389 users.
Figure 2: A question post from the OpenAI developer forum.
Step 2: Determining popularity trend.
To analyze the popularity trend of OpenAI development among developers, we first perform time series analysis. Specifically, we count the number of new posts and new users over different periods (i.e., every three months) from the forum’s inception in February 2021 until our data collection deadline in June 2024. Detailed result analysis can be found in Section 4.
Step 3: Determining difficulty level. In our primary investigation, we find that a small proportion of posts mark one of the replies as “Solution” (i.e., indicating the acceptance by the questioner). The possible reason is that users in this community do not strictly adhere to this practice, unlike those in Stack Overflow (Yazdaninia et al, 2021). Therefore, the methodology of determining difficulty levels through the proportion of accepted answers, as employed in prior studies (Wen et al, 2021; Wang et al, 2023a; Chen et al, 2020), is not applicable in our study. In our study, we mainly assess the difficulty level of questions solely based on the number of replies received. Our study assumes that a lower number of replies indicates a higher level of difficulty in addressing the question. Detailed result analysis can be found in Section 5.
Step 4: Constructing the taxonomy of challenges. Due to the high cost of manually analyzing all the crawled posts, we randomly select a statistically significant sample ensuring a 99% confidence level with a $\pm$2.5confidence interval from 29,057 questions (Aghajani et al, 2019) and remove posts that have already been deleted by OpenAI. As a result, 2,364 questions constitute the taxonomy dataset of our study. We construct the taxonomy as follows.
First, we randomly select 70% of the 2,364 posts to construct the initial taxonomy of challenges by following the previous study (Wen et al, 2021). This process involves two annotators working collaboratively. We use the open coding procedure (Seaman, 1999) to analyze the sampled posts, summarize the preliminary taxonomy, and construct a multi-level hierarchical taxonomy of challenges faced by OpenAI developers. Each annotator has two years of experience in OpenAI development and three years of experience in developing large language models (LLMs). They carefully read the questions of the sampled posts to understand the specific challenges faced by OpenAI developers and the corresponding replies. The open coding procedure is conducted as follows: the two annotators read all the sampled posts, with each post being read at least twice. During the analysis, they examine all the data contained in each post, including titles, problem descriptions, code snippets, and replies, summarizing the specific challenges discussed. Then they summarize the specific challenges discussed in the posts. Specifically, some posts describe issues encountered during the use of ChatGPT, for example, “I can’t login to ChatGPT - what should I do?”  [^5] https://community.openai.com/t/188817, which are not closely related to developers.
In addition, there are also some questions related to account problems, tips for users, etc. These posts are labeled as “Unrelated”. Since the OpenAI developer forum is a platform for all OpenAI users to communicate, some posts are official announcements, document introductions, project discussions, and project sharing. For example, “I have an idea about creating a mix of emoji”  [^6] https://community.openai.com/t/22180, these posts are temporarily labeled as “Unrelated”. If a post is labeled as “Unrelated”, it means that these posts are not considered when constructing the taxonomy because they are not closely related to OpenAI developers. For the remaining posts, the annotators use simple phrases to summarize the developer’s challenges during the first reading. Specifically, some problem descriptions are about developers consulting on parameter settings when calling OpenAI’s APIs, for example, “GPT-4-Vision-Preview fidelity/detail-parameter”  [^7] https://community.openai.com/t/477563, the annotators find that this problem is asked because developers do not know how to set API parameters. Therefore, for these issues, the annotators can temporarily describe them as “API Usage Issue, Parameter Setting”.
The two annotators continue to process all sampled posts according to this step and construct the taxonomy of challenges faced by OpenAI developers. The grouping process is iterative, with continuous modifications to the taxonomy based on the posts and optimization of the post descriptions to improve taxonomy quality. If a post involves multiple categories, it is assigned to all relevant categories. When the two annotators have different opinions on category determination, a third arbitrator is asked to coordinate conflicts. The third arbitrator has five years of software project development experience and three years of experience in developing large language models. After rigorous processing, all participants reach a consensus, and the challenges of all posts are classified, resulting in a preliminary taxonomy of challenges faced by OpenAI developers.
Finally, we refine the taxonomy of challenges and conduct a reliability analysis. Specifically, the two annotators independently label the remaining 30% of the posts based on the generated preliminary taxonomy. They label each post as “Unrelated” or assign it to the corresponding category. If some posts can not be categorized, they are temporarily labeled as “Pending”. We use Cohen’s Kappa (k(Cohen, 1960) to measure the consistency between the two annotators during the independent labeling process. The computed $k$ value is 0.812, indicating that the two annotators are almost completely agree (Landis and Koch, 1977). This demonstrates the reliability of our open coding procedure. In addition, for posts labeled as “Pending”, we introduce the third arbitrator who discusses and analyzes the posts with the two annotators to determine which category these posts ultimately belong to. If they can not be categorized into the preliminary taxonomy, new categories are added. In this phase, we incorporate nine additional leaf categories into the preliminary taxonomy framework. Finally, our constructed taxonomy covers all posts related to OpenAI development, and all participants agree on the classification. Detailed result analysis can be found in Section 6.
4 RQ1: Popularity Trend Analysis
Fig. 3 illustrates the rising trend in OpenAI’s development popularity, as evidenced by the increasing number of posts on the OpenAI forum and the surge in registered users. The figure begins at the forum’s inception, tracking the growth in new posts and users every three months. Given that the user number significantly exceeds the post number, we apply a logarithmic transformation to the actual numbers for a clearer representation of the results. This visualization reveals a significant increase in developer interest in OpenAI since 2021, underscoring the timeliness and urgency of our empirical study.
Figure 3: The number of new posts and users every three months since creating the OpenAI forum. Notice we apply a logarithmic transformation to the actual numbers for a clearer representation of the results.
Before November 2022, the growth trend was fairly consistent. However, the introduction of GPT-3.5 and its fine-tuned product, ChatGPT, to the public by OpenAI in late November 2022, marked a substantial shift. The number of posts increased by 292%, and the number of users rose by 1,302% compared to the preceding period. The release of GPT-4 to paying customers in March 2023 further accelerated this trend, with a 323% increase in posts and a 628% increase in users. This vigorous growth has persisted to the present.
Finding 1. The introduction of GPT-3.5 and ChatGPT in November 2022, followed by GPT-4 in March 2023, significantly accelerated the growth of the OpenAI community, markedly increasing the number of posts and users on the OpenAI forum.
5 RQ2: Difficulty Level Analysis
Fig. 4 illustrates the proportion of questions with varying numbers of replies. The figure shows that 54% of questions receive fewer than three replies, while only 10% receive more than ten replies. This suggests that questions related to OpenAI development are generally challenging to resolve.
Figure 4: The proportion of questions with different numbers of replies.
We summarize potential reasons as follows. First, the new and complex technological domains of OpenAI require deep expertise and skills from developers. This scarcity of qualified professionals results in limited responses. Second, the rapid evolution of OpenAI’s technology often leaves issues unresolved, leading to fewer available answers. Finally, the lack of comprehensive documentation and tailored support resources makes it difficult to address diverse developer needs, prolonging the resolution process for many questions.
Finding 2. Problems related to OpenAI development often face challenges in receiving immediate and sufficient replies.
6 RQ3: Challenge Taxonomy Construction
Fig. 5 illustrates the hierarchical taxonomy of challenges faced by OpenAI developers. In this figure, the nodes are shaded in descending levels of grey based on their depth in the hierarchy (e.g., leaf nodes are white). Each leaf node represents a specific category, while its parent node is an inner node composed of multiple subcategories. For example, API (B) is an inner category that can be further classified into four leaf categories: Faults in API (B.1), Error Messages in API Calling (B.2), and API Usage (B.3).
The proportion of posts for each category is shown in parentheses. In summary, our taxonomy includes six inner categories and 27 leaf categories. Based on this taxonomy, we find that OpenAI developers encounter a wide variety of problems across different aspects of development, highlighting the diversity of challenges in OpenAI development. Next, we describe and provide examples for each category and summarize our findings and implications for developers and the OpenAI organization.
Figure 5: Our Constructed Challenge Taxonomy for OpenAI Developers.
Finding 3. The challenges faced by OpenAI developers are multifaceted and diverse, encompassing 27 distinct categories. These range from Conceptual Questions to API Usage, from Prompt Design to Text Generation, and from Rate Limitation to Regulation.
6.1 General Questions
General questions encompass challenges that lack specific implementation details and are often posed by developers seeking foundational knowledge in OpenAI development. Our findings indicate that 26.3% of the challenges fall into this category. The detailed subcategories are discussed as follows.
Integration with Custom Applications (A.1). With the launch of ChatGPT and the availability of OpenAI’s APIs, more developers are leveraging these GPTs/APIs to build custom applications, such as translation systems and text classification systems. When integrating the API into their applications, developers often face complex and diverse technical questions. For example, developers seek to understand how to integrate custom GPTs with web page interfaces [^8] https://community.openai.com/t/589076 or inquire about the feasibility of implementing specific functionalities through API calls, such as considering specialized terminology that varies across different languages [^9] https://community.openai.com/t/584762, as shown in Fig. 6. Additionally, some developers raise concerns about the reproducibility of results in commercial text classification systems using GPT-3.5-turbo, especially after model updates [^10] https://community.openai.com/t/281658. Our findings indicate that challenges in this subcategory account for a significant 17.0%, highlighting the diverse and complex technical issues developers face when building their custom applications.
Figure 6: A sample post in the integration with custom application (A.1)
Conceptional Questions (A.2). This category encompasses questions about basic concepts or background knowledge of OpenAI development, such as how to calculate the tokens needed for each question [^11] https://community.openai.com/t/81018, whether the GPT-4 API has access to the Internet [^12] https://community.openai.com/t/468615, or whether plugins can be used with the API [^13] https://community.openai.com/t/186333. Additionally, developers inquire about the internal implementation details, such as mechanisms of the content filter [^14] https://community.openai.com/t/1095. In our taxonomy, the posts in this subcategory account for 6.4%.
Feature Suggestions (A.3). Developers often suggest adding new features related to API operations or express a desire to develop plugins for specific functionalities. For example, developers want to provide a way to view historical usage and spending by individual API key for better cost management of deployed chatbot apps using OpenAI’s GPT turbo API [^15] https://community.openai.com/t/305606. Additionally, they recommend adding search and categorization features to the plugin page, which can quickly find the desired plugins [^16] https://community.openai.com/t/215754. Challenges related to these suggestions account for 2.9% in our taxonomy.
Finding 4. The majority of the total challenges, 26.3%, fall under the category of General Questions. Within this category, the subcategory Integration with Custom Applications represents the largest proportion, accounting for 64.6% of General Questions.
Discussion and Implication: General questions often reflect fundamental challenges developers face when working with OpenAI technologies. These challenges span various aspects (such as application integration, conceptual questions, and feature suggestions). When leveraging OpenAI’s APIs to build their applications (such as translation systems, text analysis systems, speech recognition systems, and others), developers often encounter complex technical questions. These include API integration methods, performance issues, output reproducibility issues, interpretability of output content, and so on. These challenges highlight the diverse and intricate nature of API integration and the need for clearer guidelines and examples from OpenAI to assist developers in these areas. As integration with custom applications is a major challenge, OpenAI could develop dedicated resources and support mechanisms to streamline this process. This may include detailed tutorials and integration best practices. Questions in the subcategory of conceptual questions involve basic concepts and background knowledge necessary for understanding OpenAI’s technologies. Developers often inquire about token calculations, API capabilities and internal mechanisms, model versions, and other related issues. These challenges underscore the need for improved foundational documentation. To mitigate these challenges, improving educational resources will help developers grasp fundamental concepts more effectively, reducing the frequency of basic questions and accelerating their development process. For the last subcategory, developers frequently suggest new features or improvements related to API operations, such as better cost management tools, enhanced plugin and GPTs recommendation features, and support for a wide range of data formats. This subcategory reflects a proactive interest from developers in enhancing OpenAI’s tools. Therefore, OpenAI should establish a structured feedback mechanism to capture and address feature suggestions from developers. This will help prioritize feature development based on actual user needs and enhance overall user satisfaction.
6.2 API
This category contains challenges related to the OpenAI APIs. In our taxonomy, 22.9% of the total challenges fall under this category. The detailed subcategories are discussed as follows.
Faults in API (B.1). When calling OpenAI APIs, developers frequently encounter a variety of issues, such as low-quality generated content, limitations in model comprehension, text coherence problems. These issues often result in outcomes that do not meet developers’ expectations. The majority of these issues are related to unsatisfactory output, such as the presence of extraneous information (like spaces and newlines) in the API’s responses [^17] https://community.openai.com/t/578701, as well as phrase repetition in answers [^18] https://community.openai.com/t/54737. Beyond these common faults, developers face other types of faults, such as APIs failing to respond due to requests being too frequent [^19] https://community.openai.com/t/483090. Additionally, when dealing with specific technical problems, such as converting natural language into SQL queries, they might produce incorrect answers [^20] https://community.openai.com/t/426293. These examples illustrate the diverse technical challenges that can arise. This subcategory represents 8.7% of the total challenges.
Error Messages in API Calling (B.2). This category primarily includes various error messages that developers encounter when calling OpenAI APIs. For instance, developers might face request errors [^21] https://community.openai.com/t/154711 and data value capacity limit errors when calling an API for image editing [^22] https://community.openai.com/t/464513. As shown in Fig. 7, developers may also face issues related to the OpenAI server, such as gateway timeout errors [^23] https://community.openai.com/t/219369. This subcategory represents 7.5% of the total challenges.
Figure 7: A sample post in the error messages in API calling (B.2)
API Usage (B.3). Parameter configuration is a principal concern among developers, necessitating the parameter value adjustment to meet specific requirements and generate expected outputs from the API. A significant portion of these problems center on determining the appropriate values for API parameters [^24] https://community.openai.com/t/477563 (as shown in Fig. 8) and understanding the internal working and implementation mechanism (such as the response_model parameter [^25] https://community.openai.com/t/511738). Additionally, some challenges are closely associated with GPT models. For example, setting identical values for the seed parameter does not result in consistent outputs for specific GPT models [^26] https://community.openai.com/t/487245. This subcategory represents 6.7% of the total challenges.
Figure 8: A sample post in the API usage (B.3)
Finding 5. Challenges related to the API rank second, accounting for 22.9% of the total challenges. Within this category, Faults in API and Error Messages in API Calling represent 38.0% and 32.8% of the challenges, respectively.
Discussion and Implication: Our results show that developers frequently encounter various issues when using OpenAI APIs. Some of these problems are related to the specific GPT models being used. For instance, certain APIs only support the gpt-4 turbo model for specific fields like response_format, indicating potential compatibility issues. Therefore, OpenAI should ensure that their documentation clearly outlines the compatibility of different models and fields, and provide detailed examples and best practices for developers to follow. This can help mitigate confusion and reduce the frequency of these issues. Developers also encounter various error messages when calling OpenAI API. To address these issues, OpenAI should consider providing more detailed error messages that help developers understand the cause of the errors and how to fix them. Additionally, offering guidance on common parameter configurations and their implications can be beneficial. For example, addressing phrase repetition issues can be managed by adjusting the frequency_penalty parameter, where setting positive values can penalize new tokens based on their existing frequency in the text so far, thereby decreasing the model’s likelihood of repeating the same line verbatim. Parameter configuration is a principal concern among developers. To help developers, OpenAI should provide comprehensive documentation that includes detailed explanations of parameter values and their effects. Offering sample configurations and use-case scenarios can guide developers in making informed decisions about parameter settings. For example, explaining how different values for the temperature parameter affect the creativity and randomness of the generated content can help developers tailor the API output to their needs. In summary, OpenAI should ensure that the API documentation is thorough, up-to-date, and includes detailed examples and best practices. Moreover, providing more informative error messages can help developers quickly identify and resolve issues. Finally, OpenAI should monitor community forums and address recurring issues by updating the API and documentation accordingly.
6.3 Generation and Understanding
Due to powerful generation and understanding capabilities, developers can use OpenAI models to perform a variety of tasks (such as text processing, image processing, and audio processing) or fine-tune the OpenAI models for downstream tasks. the challenges in these tasks often overlap with other challenge categories, so a question (such as parameter setting issues related to image generation [^27] https://community.openai.com/t/359438) may be classified into multiple categories. Challenges related to this category account for 19.9% of the total challenges. The detailed subcategories are discussed as follows.
Text Processing (C.1). As OpenAI’s core capability, text generation helps developers to create coherent and semantically sound textual content based on the input text (i.e., the prompt). This feature allows developers to automatically generate various types of text (such as articles, stories, and dialogues). By modifying the input prompts, developers can control the style, content, and length of the generated text to meet diverse application requirements across different scenarios. This category contains challenges associated with text generation, such as encountering errors when using the text generation API [^28] https://community.openai.com/t/486260 or getting repeated responses from the davinci model [^29] https://community.openai.com/t/254241. Moreover, developers want to provide feedback on the given output to improve the upcoming outputs when using the gpt-3.5-turbo model [^30] https://community.openai.com/t/452220. This subcategory constitutes 6.8% of all challenges.
Fine-tuning GPT Models (C.2). OpenAI developers can fine-tune the models provided by OpenAI by using the training datasets of the downstream tasks. Due to the widespread use of fine-tuning, developers frequently face questions related to this functionality. For example, they often inquire about how to set the parameter n_epochs [^31] https://community.openai.com/t/339246 or request the fine-tuned model to output results in JSON format [^32] https://community.openai.com/t/525169. Additionally, developers are interested in whether they can iteratively train a fine-tuned model [^33] https://community.openai.com/t/18248. This subcategory of questions accounts for 6.7% of the total challenges.
Image Processing (C.3). Developers can influence image generation by inputting prompts into the model, allowing for the creation of images with customized features and styles. This versatile feature is useful across diverse applications, including image editing and artistic creation, offering developers opportunities for creative expression and design. However, developers often encounter challenges such as request errors when seeking image generation solutions [^34] https://community.openai.com/t/586006 or perceiving that the quality of the generated images is lower than expected [^35] https://community.openai.com/t/435187. Additionally, they are concerned about whether the generated images can be used freely without violating intellectual property rights [^36] https://community.openai.com/t/23140. This subcategory constitutes 2.5% of all challenges.
Embedding Generation (C.4). By using the Embedding APIs provided by OpenAI, developers can leverage powerful text representation capabilities to convert text into high-dimensional vector representations. These vectors can capture the semantics and contextual information of the text, which can be used for tasks such as semantic search and text clustering. For example, developers may encounter PermissionError when creating embeddings [^37] https://community.openai.com/t/587579. Additionally, they often inquire whether modifying user queries can improve the accuracy of semantic search [^38] https://community.openai.com/t/393047 and how to determine the length of embedded content [^39] https://community.openai.com/t/111471. Challenges related to this subcategory account for 1.8% of the total challenges.
Audio Processing (C.5). Processing audio is one of the key functionalities provided by OpenAI. Developers can utilize advanced speech models (such as Whisper) to perform transformations, analysis, and generation of audio data, thereby supporting applications such as speech recognition and audio generation. For instance, developers might encounter random text in the output after the API recognizes audio [^40] https://community.openai.com/t/287544 or face InvalidRequestError when calling the whisper-1 API [^41] https://community.openai.com/t/433315, as shown in Fig. 9. Additionally, they often ask how to combine audio processing with image processing [^42] https://community.openai.com/t/22912. Challenges in this subcategory constitute 1.4% of all challenges.
Figure 9: A sample post in audio processing (C.5).
Vision Capability (C.6). The Vision API uses pre-trained vision models to quickly and accurately identify objects, scenes, and features in images, providing developers with powerful image processing capabilities. This functionality is widely applicable in areas such as image recognition, intelligent surveillance, and medical image analysis. However, developers may encounter error messages when using the gpt-4-vision-preview API [^43] https://community.openai.com/t/546593 or find that it occasionally misinterprets information in images [^44] https://community.openai.com/t/504043. They also ask usage-related questions, such as how to add a parameter to view image resolution [^45] https://community.openai.com/t/477563. This subcategory constitutes 0.7% of all challenges.
Finding 7. The category of Generation and Understanding, the third-largest category, has 34.2% of challenges concentrated in Text Processing and accounts for 19.9% of all challenges.
Discussion and Implication: The analysis of the various subcategories within the ”Generation and Understanding” category reveals several insights into the challenges developers face and the implications for improving OpenAI’s models and their usage. (1) API Usage Issues. A significant portion of the challenges are related to the practical use of APIs. Issues such as repeated responses from the Davinci model, errors in embedding API usage, and problems with the Whisper API during audio processing are common. Developers often struggle with these API-related problems, seeking solutions and best practices for effective implementation. Therefore, detailed, updated documentation should be provided, especially for frequently encountered issues. For instance, recommending the use of the latest APIs, such as the Chat Completions API instead of older models like Davinci, can help mitigate repeated response issues. Providing best practices for embedding API usage, such as adding an offset vector to improve accuracy, can enhance developer experience. Similarly, sharing solutions for Whisper API errors, like addressing audio recording and encoding issues, can be beneficial. (2) Conceptual Challenges. Developers often face conceptual challenges, particularly with fine-tuning GPT models, parameter settings, iterative training, and output formatting, indicating a need for a deeper understanding of these processes. Therefore, creating comprehensive guides and tutorials on fine-tuning, common parameter settings, and iterative training techniques, can address many of these questions. (3) Application-Specific Issues. Challenges also arise from specific applications. Taking image processing as an example, developers encounter issues like translating prompts from other languages (such as French) to English, intellectual property concerns with generated images, and improving image recognition accuracy by adjusting contrast or rotation. Therefore, enhancing multilingual support for image processing prompts can help generate the desired images regardless of the input language. Providing clear guidelines on the intellectual property rights of generated images can address developers’ concerns and promote more confident use of the technology. Sharing techniques for improving image recognition accuracy, such as contrast adjustment and rotation correction, can help developers achieve better results in their applications.
6.4 Non-functional Properties
The importance of non-functional properties cannot be ignored by OpenAI developers. These properties include aspects such as API call costs, rate limitations, and data security. In our taxonomy, 15.4% of the challenges belong to this category. The detailed subcategories are discussed as follows.
Costs (D.1). Challenges in this category are related to API call costs. In the context of OpenAI development, costs are determined by the number of tokens utilized in each API call, particularly for generated content. Developers express concerns regarding the calculation of token usage and the cost [^46] https://community.openai.com/t/496993. They seek strategies to reduce costs by minimizing token consumption [^47] https://community.openai.com/t/525033. Responses to these challenges typically involve providing documentation-related information or offering strategies that consider factors such as the chosen model, input content length, and response content length. This category constitutes 3.6% of all challenges.
Rate Limitation (D.2). Rate limitation is typically implemented to safeguard the stability and reliability of OpenAI services, preventing service interruptions or performance degradation due to excessive API usage. Developers need to understand the rate-limiting rules of the API, including the maximum number of calls per second, per minute, or per hour for different versions of the API, and how to manage call frequency in a sensible way to avoid triggering rate limits and causing service interruptions or exceptions. For example, developers encounter RateLimitError when calling gpt-3.5-turbo-0301484848https://community.openai.com/t/566696. Additionally, developers inquire whether rate limits are shared among different APIs [^49] https://community.openai.com/t/360331. Furthermore, developers ask about methods to increase the API rate limits [^50] https://community.openai.com/t/248374. These types of questions account for 3.2% of the total challenges.
Regulation (D.3). When utilizing the OpenAI APIs, developers should strictly adhere to relevant regulations. For instance, OpenAI provides guidelines for the release of specific models or features and determines which users are granted permission to utilize them. Developers should review and abide by the terms of service and usage policies provided by OpenAI to ensure compliance with commercial and copyright regulations during API calls. For example, developers want to know whether OpenAI API can be used for commercial purposes [^51] https://community.openai.com/t/193738, or whether the images generated by DALL-E can be sold [^52] https://community.openai.com/t/33538. Additionally, developers encounter issues where prompts violate content policy rules [^53] https://community.openai.com/t/468967. This category constitutes 3.0% of all challenges.
Promotion (D.4). Developers leverage forums as platforms to promote their developed plugins and GPTs, aiming to gather user feedback and potentially generate revenue. To this end, they offer comprehensive introductions to their creations, innovative functionalities, and usage guidelines. For instance, a developer promotes a plugin named “ShellMaster” [^54] https://community.openai.com/t/285122. Additionally, developers complain that the current GPT store lacks discoverability, meaning that many excellent GPTs are likely unknown to most people. They hope to add recommendations for users [^55] https://community.openai.com/t/585652. This category represents 2.1% of the challenges.
Token Limitations (D.5). In addition to rate limiting, OpenAI imposes restrictions on the number of tokens for input and output during API calls to prevent abuse of API resources. Therefore, developers should understand the allocation, usage rules, and limitations of API tokens, as well as how to manage and optimize their usage. Specifically, developers find that the context length of the API is not fully utilized [^56] https://community.openai.com/t/291384. Additionally, developers seek methods to address token limitations  [^57] https://community.openai.com/t/4914. Furthermore, they inquire about solutions for handling large-scale datasets that exceed the token limitation and ineffective data splitting when using the gpt-4-1106-preview model [^58] https://community.openai.com/t/566336, as shown in Fig. 10. This category constitutes 2.0% of the challenges.
Figure 10: A sample post in token limitations (D.5)
Security (D.6). Ensuring security and preventing data leakage during API calls are the main concerns for OpenAI developers. They should prioritize protecting the privacy and integrity of data, minimizing the risk of data leaks due to security vulnerabilities or improper configurations. For example, developers express apprehension about potential data security risks when utilizing the API to process their uploaded files [^59] https://community.openai.com/t/354485. Additionally, they want to know whether using the API within internal networks can lead to data leakage [^60] https://community.openai.com/t/75725. This category constitutes 0.8% of all challenges.
Privacy (D.7). Developers are advised to carefully review OpenAI’s privacy policy to gain insights into the data collection and processing practices during API calls and to understand the measures in place to protect users’ personal privacy. Furthermore, since API keys serve as critical qualifications for developer authentication and authorization, it is imperative for developers to prioritize the secure storage and management of these keys to prevent leaks or misuse. For example, developers create a desktop application based on ChatGPT and want to know if each user can use their API key within the program for billing purposes [^61] https://community.openai.com/t/161185. Additionally, developers have concerns regarding potential privacy issues associated with GPTs [^62] https://community.openai.com/t/496343. This subcategory constitutes 0.7% of all challenges.
Finding 8. 15.4% of the total challenges are in Non-functional Properties, covering seven subcategories. Cost and Rate Limitation are the top two categories, together accounting for 44.2% of the challenges in Non-functional Properties.
Discussion and Implication: Challenges such as API call costs, rate limitations, and token limitations are tightly linked to the development and usage of OpenAI’s services. Developers often express concerns about the costs associated with API calls, which are influenced by the choice of model and the number of tokens used in each request. Similarly, rate limitations are put in place to ensure service stability, but developers need to understand these limits and manage their API call frequencies accordingly. Token limitations pose additional challenges, especially when developers need to handle large datasets or require extensive context in their applications. Therefore, OpenAI should develop and provide tools that help developers accurately calculate and manage their token usage. This includes offering cost optimization strategies tailored to different models and usage scenarios. Moreover, providing detailed guidelines on model selection, including trade-offs between cost and performance, will help developers make informed decisions that align with their budget constraints.
General challenges such as safety and privacy are paramount when using AI services. Developers should ensure that their applications adhere to safety standards and protect user privacy. Therefore, OpenAI should continue to promote the use of its free moderation API, which helps reduce the frequency of unsafe content in completions. This tool should be integrated into the development process to automatically flag and filter potentially harmful outputs. Moreover, in high-stakes domains and code generation, it is crucial to have a human review of outputs before they are used in practice. This ensures that any limitations of the system are accounted for, and provides a safety net to verify the correctness and appropriateness of the generated content. Finally, limiting the amount of text a user can input into prompts and restricting the number of output tokens can help mitigate the risks of prompt injection and reduce the chances of misuse. These limitations should be implemented as part of best practices for safe AI deployment.
6.5 GPT Builder
To meet specific user needs, OpenAI initially introduced the ChatGPT plugin, allowing developers to design plugins with various functionalities based on user requirements. Recently, OpenAI launched GPTs with similar functionality. However, due to their later release, there are fewer posts related to GPTs. Although OpenAI has deprecated plugins and recommends using GPTs as a replacement for users’ personalized needs, the OpenAI developer forum still lists plugins as an important category and combines plugins and GPTs into a single category. Moreover, plugin development and GPT development share many similarities in terms of development goals, technical foundations, and user usage. Therefore, we refer to the categories on the forum and collectively call both ChatGPT plugins and GPTs “GPT Builders” and analyze the related challenges.
Development (E.1). In plugin development, developers encounter various challenges, including general and technical problems. Among general problems, developers seek guidance on selecting the most convenient IDE for plugin development [^63] https://community.openai.com/t/404743 and whether Markdown code can be integrated into plugin development [^64] https://community.openai.com/t/199399. Another category of challenges is related to technical issues. For instance, developers encounter a FatalServerError while developing a plugin [^65] https://community.openai.com/t/297501 or face parsing errors when attempting to connect GPTs to a cloud website [^66] https://community.openai.com/t/545528, as shown in Fig. 11. Additionally, developers ask how to use multiple GPTs simultaneously to handle the same task [^67] https://community.openai.com/t/552649. These challenges account for 11.2% of the total challenges.
Figure 11: A sample post in development (E.1)
Testing (E.2). Following the development of plugins or GPTs, developers engage in software testing to ensure the correctness of the functionality implementation. During this process, developers encounter various issues. For instance, the openapi.yaml file might fail initial validation because it makes an HTTPS request instead of HTTP when testing the plugin on a local server [^68] https://community.openai.com/t/254680, or developers encounter errors while testing if the plugin is working correctly [^69] https://community.openai.com/t/166239. These challenges represent 0.9% of the total challenges.
Finding 9. GPT Builder accounts for 12.1% of the total challenges. Development dominates this category, accounting for 92.6% of the challenges in GPT Builder.
Discussion and Implication: Developers face numerous general challenges when developing plugins and GPTs, particularly in selecting the appropriate development environments and tools. This involves decisions about which Integrated Development Environments (IDEs) to use, which programming languages best suit their project needs, and how to set up their development environments efficiently. Therefore, OpenAI should develop comprehensive guides that cover the entire development lifecycle of plugins and GPTs, including recommendations for IDEs, programming languages, and other essential tools. This will help developers make informed choices that enhance their productivity and the quality of their work. Moreover, providing clear examples and templates can streamline the development process, particularly for those at varying skill levels. These resources can include step-by-step tutorials, sample projects, and best practices for developing and debugging, ensuring that developers have the support they need from project initiation to completion.
After the development phase, rigorous testing is crucial to ensure the correct functioning of plugins and GPTs. Therefore, OpenAI should provide comprehensive automated testing tools that allow developers to test their plugins and GPTs efficiently. Automated testing can help identify bugs and issues early in the development cycle, reducing the time spent on manual testing and allowing for more consistent and repeatable tests. Moreover, these tools should support various testing methodologies, including unit tests, integration tests, and acceptance tests. By covering a broad range of scenarios, developers can ensure their applications are robust and reliable.
6.6 Prompt
Prompts play a pivotal role in OpenAI API interactions, offering developers a mechanism to enhance the quality of the content produced. This category primarily focuses on challenges related to prompts, such as how to design and optimize them. Our findings indicate that 3.4% of the challenges fall into this category. Below, we analyze the specific subcategories it contains.
Prompt Design (F.1). Developers often struggle with how to provide specific prompts to achieve the desired results and how to adjust their existing prompts. This includes techniques like targeting specific topics and managing the length of generated content to guide the model more accurately. For example, they seek advice on prompts to regenerate part of the output [^70] https://community.openai.com/t/214339(as shown in Fig. 12) and how to adjust prompts to improve the conversational style with the GPT model [^71] https://community.openai.com/t/301950. Challenges related to prompt design constitute 2.3% of the total questions.
Figure 12: A sample post in prompt design (F.1)
Retrieval Augmented Generation (F.2). Retrieval Augmented Generation (RAG) Lewis et al (2020) is a technique that enhances the quality and relevance of generated text by integrating information retrieval with generation processes. This technique combines the strengths of retrieval and generation, enabling the model to better grasp the context and produce text that meets user expectations. Developers encounter issues where RAG fails to improve output quality [^72] https://community.openai.com/t/550286 or cause hallucinations after processing [^73] https://community.openai.com/t/408275. Questions also arise about preventing the RAG system from answering unrelated questions [^74] https://community.openai.com/t/434871. Challenges related to RAG account for 0.4% of the total challenges.
Chain of Thought (F.3). The Chain of Thought (CoT) technique Wei et al (2022) is crucial for maintaining coherence and logical progression during text generation. It requires developers to craft input prompts that encourage the generation of text following a logical and semantic chain of thought. Achieving this involves careful selection of keywords, phrases, and sentence structures to steer the model toward producing coherent and structured text. Developers need a robust understanding of language, logical reasoning, and text generation capabilities, along with the skill to evaluate the model’s behavior and outputs. Ensuring semantic coherence and meeting expectations might prompt questions on effectively utilizing the CoT technique to construct prompts [^75] https://community.openai.com/t/128180 and seeking best practices for CoT in specific scenarios [^76] https://community.openai.com/t/17367. Challenges associated with this technique account for 0.2% of the total challenges.
In-context Learning (F.4). In-context learning (ICL) (Dong et al, 2022) involves the model’s ability to learn and interpret based on the contextual information supplied within the input prompt. By meticulously designing these prompts, developers can embed specific contextual clues, enabling the model to deduce and produce content reflective of that context. For instance, there may be situations where, despite being provided with examples, the model repetitively generates identical outputs [^77] https://community.openai.com/t/17866. Developers seek advice on the recommended approach to provide context to follow-up questions to retrieve relevant context for completion [^78] https://community.openai.com/t/280666. Challenges related to in-context learning represent 0.2% of the total challenges.
Zero-shot Prompting (F.5). Zero-shot prompting Kojima et al (2022) allows developers to leverage the model’s capabilities without needing sample data. Common issues include comparing the efficacy of zero-shot prompting versus fine-tuning for specific tasks [^79] https://community.openai.com/t/289714. Recently, some developers observed that zero-shot prompting works as expected on the OpenAI Playground but fails to generate valid responses when called via the Python API [^80] https://community.openai.com/t/216108. Challenges related to zero-shot prompting constitute 0.2% of the total questions.
Tree of Thoughts (F.6). Tree of Thoughts (ToT) (Yao et al, 2024) is an advanced structured Chain of Thought (CoT) technique that guides the model toward generating text that is hierarchically organized and logically coherent. By creating detailed input prompts that integrate information from various branches and nodes within the ToT framework, developers enable the model to produce outputs with a clear hierarchy and logical structure. This technique assists the model in adhering to the ToT’s structure, facilitating the creation of more organized and logically structured outputs. Researchers have found that ToT significantly enhances language models’ problem-solving abilities in tasks requiring non-trivial planning or search, such as the Game of 24, creative writing, and mini crosswords [^81] https://community.openai.com/t/226512. Challenges related to ToT account for 0.1% of the total challenges.
Finding 6. 3.4% of the total challenges are related to Prompt. Prompt Design constitutes 67.6% of this category, which is the highest proportion among the six subcategories.
Discussion and Implication: Developers often encounter challenges when crafting prompts that achieve specific outcomes, manage content length, and maintain a conversational style. Prompt engineering involves feeding the model useful phrases and instructions to elicit helpful responses, which includes clearly stating the desired topic or task, and specifying constraints to guide responses within defined boundaries. Therefore, OpenAI should provide comprehensive educational resources on prompt engineering, including tutorials, webinars, and documentation. These resources should cover the principles of effective prompt design, common pitfalls, and best practices for crafting prompts that yield accurate and useful responses. Moreover, developing interactive tools that allow developers to experiment with and optimize prompts in real-time would be highly beneficial. Finally, establishing a community platforms where developers can share successful prompts, discuss challenges, and collaborate on solving common issues is crucial.
Recent advancements in prompt design (such as In-context Learning, Chain-of-Thought) offer powerful techniques for enhancing the performance of OpenAI models. These techniques involve using contextual information, reasoning through chains of logic, and augmenting responses with external data to improve the quality and relevance of the output. However, in our analysis, the number of posts discussing these advanced techniques is relatively low, indicating that developers are not paying enough attention to these advanced technologies. Therefore, OpenAI should offer extensive documentation and resources on these advanced techniques. This documentation should include detailed explanations of the underlying principles, common pitfalls, and practical examples to help developers effectively implement these techniques. Providing step-by-step guides and use case scenarios can make these resources more accessible and practical. Moreover, providing interactive tools that allow developers to experiment with and optimize these advanced techniques in real-time would be invaluable.
7 Threats to Validity
In this section, we discuss the threats to the validity of our empirical study, ensuring a comprehensive understanding of potential limitations.
7.1 Internal Validity
Expertise in Constructing the Challenge Taxonomy: One internal threat is related to the expertise involved in constructing the challenge taxonomy. To mitigate this, individuals involved in the taxonomy construction are required to have relevant experience in OpenAI development and a foundational understanding of large language models. This ensures that the taxonomy accurately reflects the challenges faced by developers.
Determining Question Difficulty Levels: Another internal threat is the method used to determine the difficulty level of questions posed on the OpenAI forum. Our analysis shows that only 7.0% of the sampled posts had one reply marked as a “Solution”, Hence, the difficulty level is primarily determined by the number of replies each question receives rather than relying on accepted answers.
Manual Labeling of Topics: We manually label the topics, which can introduce bias. To minimize this threat, two different authors independently label the topics, and a third author, a domain expert, validates these labels. Any conflicts are resolved through discussions among the authors, thereby reducing labeling bias to an acceptable minimum.
Selection of Forum Posts: Our study selectively focuses on posts within specific categories like API, Prompting, ChatGPT, and GPT builders, excluding others such as Community, Announcements, Documentation, and Forum feedback. This selection is based on the observation that excluded categories rarely discuss challenges related to OpenAI development.
7.2 External Validity
Scope of Data: An external challenge arises from the temporal limitation of our data, encompassing posts up to June 2024. As OpenAI continues to develop new applications, new challenges will likely emerge, which are not captured in our current empirical study.
Incorrect initial categorization: Some developers may choose the wrong category when posting. For example, the issue is actually related to API calls, but the developer categorizes it under Prompt. To address this issue, when categorizing posts, we carefully read all the information (including questions and replies) and analyze the correct category to mitigate such errors.
7.3 Construct Validity
Accuracy of Taxonomy Categories: There is a risk that the taxonomy categories may not fully capture all relevant challenges faced by OpenAI developers. To address this, we iteratively refine the taxonomy based on feedback and consensus among experienced annotators, ensuring a comprehensive representation of challenges.
8 Related Work
8.1 OpenAI Models for Software Engineering Tasks
Recently, OpenAI models (such as GPT-3.5, GPT-4, and GPT-4o) has been successfully applied to assist various software engineering tasks, demonstrating its versatility and effectiveness in the field. Due to the limited length of the paper, we only listed some relevant studies on popular software engineering tasks.
One significant application is in code generation, where OpenAI models have been used to refine and improve code creation processes, making it easier for developers to produce high-quality code efficiently (Liu et al, 2023b, a). In program repair, OpenAI models have shown promise by automatically identifying and fixing bugs, thereby reducing the time and effort required for manual debugging (Xia and Zhang, 2023). Moreover, ChatGPT has been utilized in bug report management, streamlining the process of handling and categorizing bug reports, which enhances the overall efficiency of software maintenance (Zhang et al, 2023). In app testing, OpenAI models aid in automating the testing process, ensuring more comprehensive test coverage and faster identification of issues (Liu et al, 2023c). Additionally, OpenAI models has been employed in vulnerability detection, where it helps in identifying potential security vulnerabilities in code, contributing to the development of more secure software applications (Fu et al, 2023; Chen et al, 2023; Lu et al, 2024). Lastly, for source code summarization, ChatGPT has been used to generate concise and accurate summaries of source code, aiding developers in understanding and documenting code more effectively (Sun et al, 2023; Zhao et al, 2023a).
8.2 Studies on Developers’ Challenges
In traditional software development, developers face a broad range of challenges for different application domains (such as mobile applications (Rosen and Shihab, 2016), web applications (Scoccia et al, 2021), and concurrency development (Ahmed and Bagherzadeh, 2018)). These include understanding complex requirements, managing project scope, ensuring code quality and maintainability, handling security vulnerabilities, and integrating various software components.
In the realm of deep learning, developers face distinct challenges that differ from traditional software development. These include managing and preprocessing large datasets, tuning hyperparameters, ensuring model interpretability, and addressing ethical concerns related to bias and fairness. Deployment challenges (Chen et al, 2020) are also significant, involving the integration of deep learning models into production systems, handling performance bottlenecks, and ensuring compatibility with existing infrastructure. Empirical studies have highlighted specific issues such as bugs in model optimization (Guan et al, 2023), compatibility problems (Wang et al, 2023b), and performance-related concerns that complicate the deployment and maintenance of deep learning systems (Cao et al, 2022).
Different from previous studies, we aim to fill the gap in understanding the specific challenges faced by OpenAI developers. By systematically categorizing and analyzing the posts in the OpenAI developer forum, we provide a comprehensive overview of the challenges developers encounter. Unlike previous studies that focus on broader categories or specific domains, our research delves into the unique difficulties of OpenAI development, particularly with OpenAI’s APIs and tools. Our findings offer valuable insights for both developers and the OpenAI organization, guiding future improvements in developer support and tool development. This study not only highlights common challenges but also identifies unique challenges faced by developers, providing a foundation for targeted solutions and enhancements in OpenAI’s development ecosystem.
9 Conclusion
In this study, we conduct a comprehensive analysis of the challenges faced by OpenAI developers. By examining relevant posts on the OpenAI developer forum, we observe that OpenAI-related development is gaining significant traction, with developers encountering numerous challenging issues compared to traditional software development. Our goal is to analyze the underlying challenges reflected in these questions. We investigate the popularity trends on the OpenAI developer forum and the difficulty levels of the problems raised by developers. Subsequently, we manually inspect 2,364 sampled posts related to OpenAI development and construct a detailed taxonomy consisting of 6 main categories and 27 subcategories, representing the challenges that OpenAI developers encounter. Finally, we discuss the implications of our findings for various stakeholders, including developers, researchers, and the OpenAI organization.
For future research, there are several avenues to explore. First, expanding the analysis to include posts from other platforms, such as GitHub and Stack Overflow, can provide a more comprehensive view of the challenges faced by OpenAI developers. Second, an in-depth study of the solutions proposed in response to the identified challenges can yield valuable insights into effective strategies and best practices. Finally, automating the categorization process using machine learning techniques can improve the efficiency and accuracy of the analysis.
Acknowledgements.
Xiang Chen and Chaoyang Gao have contributed equally to this work and they are co-first authors. Xiang Chen is the corresponding author. This research was partially supported by the National Natural Science Foundation of China (Grant no. 61202006) and the Postgraduate Research & Practice Innovation Program of Jiangsu Province (Grant no. SJCX24_2022).
Data Availability Statements
We share our collected posts, analysis scripts, and taxonomy of challenges in a GitHub repository (https://github.com/judeomg/OADF) for open science.
Declarations
Conflict of Interest The authors declare that they have no conflict of interest.